# FINITE SUBGROUPS OF MAXIMAL ORDER OF THE CREMONA GROUP OVER THE RATIONALS

### v1

@@ 9. Automorphism group of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ over $\mathbb{Q}$
Let $G$ be a finite group in $\operatorname{PGL}_{2}(\mathbb{Q})$. [Bea10, Theorem 4.2.] shows that the group $G$ can be of size $12$ at most. Moreover, the $2-$groups are of order at most $2^{2}$ and the $3-$groups are of order at most $3$.
@@
[Bea10] Let $G\subset\operatorname{PGL}_{2}(\mathbb{Q})$ be a finite subgroup, then the group $H:=(G\times G)\rtimes\mathbb{Z}/2\mathbb{Z}$ is a finite subgroup of $\operatorname{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$. Furthermore, if $G\cong D_{6}$, then the group $H$ is of order $288$.
@@
10. Del Pezzo surfaces of degree $2$
A Del Pezzo surface of degree $2$ defined over $\mathbb{Q}$ is given by the blow-up of $7$ points of $\mathbb{P}^{2}$ in general position (no six of them are on a conic and no three are collinear). Let $A_{1},\ldots,A_{7}\in\mathbb{P}^{2}$ be the $7$ points and $\pi:S\rightarrow\mathbb{P}^{2}$ the blow-down. The system of cubics passing through the $7$ points gives a morphism $\eta:S\rightarrow\mathbb{P}^{2}$ whose linear system is $\mid-K_{S}\mid$ and which is a double covering ramified over a smooth quartic curve $\Gamma$ of $\mathbb{P}^{2}$. (See [BB00, Fer04, DI09]):
${S}$${\mathbb{P}^{2}}$${\mathbb{P}^{2}.}$$\scriptstyle{\pi}$$\scriptstyle{\eta}$$\scriptstyle{{C^{3}(A_{1},..,A_{7})}~{}~{}~{}~{}~{}}$
@@ 10. Del Pezzo surfaces of degree $2$
Let $S$ be a Del Pezzo surface of degree $2$ defined over $\mathbb{Q}$. By [Kol96, Theorem III.3.5] and [KSC04, Corollary 3.54], over $\mathbb{C}$, there exists a smooth quartic curve $\Gamma$ of $\mathbb{P}^{2}$ given by the equation $F(x,y,z)=0$ where $F$ is homogeneous of degree $4$ such that the surface $S$ can be viewed in the weighted projected space $\mathbb{P}(2,1,1,1)$. The morphism to $\mathbb{P}^{2}$ is the anti-canonical morphism, thus it is defined over $\mathbb{Q}$, so the quartic curve is also defined over $\mathbb{Q}$. Since every automorphism of $S$ must leave the quartic $w=F(x,y,z)=0$ invariant, we have a homomorphism $\rho:\operatorname{Aut}(S)\rightarrow\operatorname{Aut}(\mathbb{P}^{2},\Gamma)$. By [Bla06, Proposition 8.3.1], the exact sequence
@@
splits, where $\sigma$ denotes the Geiser involution $\sigma(w,x,y,z)=(-w,x,y,z)$. (See [Bla06, Lemma 8.3.2]). Furthermore, $\operatorname{Aut}_{\mathbb{C}}(S)\cong~{}\langle\sigma\rangle\times\operatorname{Aut}_{\mathbb{C}}(\mathbb{P}^{2},\Gamma)$ (see [Bla06, Proposition 8.3.1]). Since $\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2},\Gamma)\subset\operatorname{PGL}_{3}(\mathbb{Q})$, by Lemma 6.12, $\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2},\Gamma)$ is of size $24$ at most. This implies to $\mid\operatorname{Aut}_{\mathbb{Q}}(S)\mid=2\cdot\mid\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2},\Gamma)\mid\leq 2\cdot 24=48$. ∎
@@
This surface $S$ contains $3$ disjoint $(-1)-$curves, given by
@@
By contracting these three $(-1)-$curves, there is thus a contraction morphism $\eta:S\rightarrow Y$, where $Y$ is a Del Pezzo surface of degree $5$, then $Y(\mathbb{Q})\neq\phi$, so the surface $S$ is rational. Indeed, this follows from [Man66, Theorem 3.15]. Furthermore, the automorphism group $\operatorname{Aut}(S)$ is generated by
@@
11. Del Pezzo surfaces of degree $1$
A Del Pezzo surface $S$ of degree $1$ is given by the blow-up of $8$ points of $\mathbb{P}^{2}$, in general position. The linear system $\mid-2K_{S}\mid$ induces a degree $2$ morphism onto a quadric cone in $Q\subset\mathbb{P}^{2}$, ramified over the vertex $v$ of $Q$ and a smooth curve $C$ of genus $4$. Moreover, $C$ is the intersection of $Q$ with a cubic surface. (See [BB00, Fer04, DI09]).
@@ 11. Del Pezzo surfaces of degree $1$
Let $S$ be a Del Pezzo surface of degree $1$ defined over $\mathbb{Q}$. By [Kol96, Theorem III.3.5] and [KSC04, Corollary 3.54], up to a change of coordinates, we may assume that the surface $S$ has the equation
@@
$[w:x:y:z]\mapsto[\lambda\cdot w+p_{3}+z\cdot p_{1}:ax+by:cx+dy:\mu\cdot z+p_{2}],$
@@
where $g_{i}=f_{i}(ax+by,cx+dy)$. This should be a multiple of the Equation 8. The degree $1$ coefficient in $w$ is $2\cdot\lambda\cdot w\cdot(p_{3}+z\cdot p_{1})$ and needs to be zero, so $p_{3}=0$ and $p_{1}=0$. Then, the degree $2$ coefficient of $z$ is $-3\cdot\mu^{2}\cdot p_{2}$ so needs to be zero. This implies that every automorphism is of the form
$[w:x:y:z]\mapsto[\lambda\cdot w:ax+by:cx+dy:\mu\cdot z].$
@@
where $g_{i}=f_{i}(ax+by,cx+dy)$. We find that $\lambda^{2}=\mu^{3}$, so by Lemma 6.3, there exists $u\in\mathbb{Q}^{*}$ such that $\lambda=u^{3}$ and $\mu=u^{2}$. As $[w:x:y:z]=[u^{3}w:ux:uy:u^{2}z]$, we can replace $\lambda$ with $\lambda/u^{3}=1$ and $\mu$ with $\mu/u^{2}=1$. Every element is then of the form
$[w:x:y:z]\mapsto[w:ax+by:cx+dy:z],$
where $\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in\operatorname{GL}_{2}(\mathbb{Q})$. Moreover the matrix is unique, as if we multiply everything with $[s^{3}:s:s:s^{2}]$, then to obtain $s^{2}=s^{3}=1$ we need $s=1$. The group of automorphism is then a finite subgroup of $\operatorname{GL}_{2}(\mathbb{Q})$. In $\operatorname{GL}_{2}(\mathbb{Q})$, every finite subgroup divides $12$ (see [New72, Chap. IX, section 14]), then $\mid\operatorname{Aut}(S_{\mathbb{Q}})\mid\leq 12$. Replacing again in the Equation 8, we see that $f_{i}(ax+by,cx+dy)=f_{i}$ for each $i$. The group of automorphism is then just
@@
where $\alpha^{6}=\beta^{2}=id$ and $\beta\alpha\beta=\alpha^{-1}$. Hence, the group $\operatorname{Aut}(S_{\mathbb{Q}})\cong D_{6}$ and it has $12$ elements. Moreover, If $\left\{\lambda=-{\frac{1}{5}},\mu=-{\frac{6}{5}}\right\}$, this surface $S$ contains $4$ disjoint $(-1)-$curves, given by
@@
By contracting these four $(-1)-$curves, there is thus a contraction morphism $\eta:S\rightarrow Y$, where $Y$ is a Del Pezzo surface of degree $5$, then $Y(\mathbb{Q})\neq\phi$, so the surface $S$ is rational. Indeed, this follows from [Man66, Theorem 3.15].

### v2

@@ 9. Automorphism group of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ over $\mathbb{Q}$
Let $G$ be a finite group in $\operatorname{PGL}_{2}(\mathbb{Q})$. [3, Theorem 4.2.] shows that the group $G$ can be of size $12$ at most. Moreover, the 22-groups are of order at most $2^{2}$ and the 33-groups are of order at most $3$.
@@
[3] Let $G\subset\operatorname{PGL}_{2}(\mathbb{Q})$ be a finite subgroup, then the group $H:=(G\times G)\rtimes\mathbb{Z}/2\mathbb{Z}$ is a finite subgroup of $\operatorname{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$. Furthermore, if $G\cong D_{6}$, then the group $H$ is of order $288$.
@@
10. Del Pezzo surfaces of degree 22
A Del Pezzo surface of degree $2$ defined over $\mathbb{C}$ is given by the blow-up of $7$ points of $\mathbb{P}^{2}$ in general position (no six of them are on a conic and no three are collinear). Let $A_{1},\ldots,A_{7}\in\mathbb{P}^{2}$ be the $7$ points and $\pi:S\rightarrow\mathbb{P}^{2}$ the blow-down. The system of cubics passing through the $7$ points gives a morphism $\eta:S\rightarrow\mathbb{P}^{2}$ whose linear system is $\mid-K_{S}\mid$ and which is a double covering ramified over a smooth quartic curve $\Gamma$ of $\mathbb{P}^{2}$. (See [1, 7, 8]):
[…]
@@ 10. Del Pezzo surfaces of degree 22
Let $S$ be a Del Pezzo surface of degree $2$ defined over $\mathbb{Q}$. By [15, Theorem III.3.5] and [14, Corollary 3.54], over $\mathbb{C}$, there exists a smooth quartic curve $\Gamma$ of $\mathbb{P}^{2}$ given by the equation $F(x,y,z)=0$ where $F$ is homogeneous of degree $4$ such that the surface $S$ can be viewed in the weighted projected space $\mathbb{P}(2,1,1,1)$. The morphism to $\mathbb{P}^{2}$ is the anti-canonical morphism, thus it is defined over $\mathbb{Q}$, so the quartic curve is also defined over $\mathbb{Q}$. Since every automorphism of $S$ must leave the quartic $w=F(x,y,z)=0$ invariant, we have a homomorphism $\rho:\operatorname{Aut}(S)\rightarrow\operatorname{Aut}(\mathbb{P}^{2},\Gamma)$. By [4, Proposition 8.3.1], the exact sequence
@@
splits, where $\sigma$ denotes the Geiser involution $\sigma(w,x,y,z)=(-w,x,y,z)$. (See [4, Lemma 8.3.2]). Furthermore, $\operatorname{Aut}_{\mathbb{C}}(S)\cong~\langle\sigma\rangle\times\operatorname{Aut}_{\mathbb{C}}(\mathbb{P}^{2},\Gamma)$. (see [4, Proposition 8.3.1]). Since $\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2},\Gamma)\subset\operatorname{PGL}_{3}(\mathbb{Q})$, by Lemma 6.12, $\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2},\Gamma)$ is of size $24$ at most. This implies to $\mid\operatorname{Aut}_{\mathbb{Q}}(S)\mid=2\cdot\mid\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2},\Gamma)\mid\leq 2\cdot 24=48$. ∎
@@
This surface $S$ contains $3$ disjoint (1)(-1)-curves, given by
@@
By contracting these three (1)(-1)-curves, there is thus a contraction morphism $\eta:S\rightarrow Y$, where $Y$ is a Del Pezzo surface of degree $5$, then $Y(\mathbb{Q})\neq\emptyset$, so the surface $S$ is rational. Indeed, this follows from [17, Theorem 3.15]. Furthermore, the automorphism group $\operatorname{Aut}(S)$ is generated by
@@
11. Del Pezzo surfaces of degree 11
A Del Pezzo surface $S$ of degree $1$ defined over $\mathbb{C}$ is given by the blow-up of $8$ points of $\mathbb{P}^{2}$, in general position. The linear system $\mid-2K_{S}\mid$ induces a degree $2$ morphism onto a quadric cone in $Q\subset\mathbb{P}^{2}$, ramified over the vertex $v$ of $Q$ and a smooth curve $C$ of genus $4$. Moreover, $C$ is the intersection of $Q$ with a cubic surface. (See [1, 7, 8]).
@@ 11. Del Pezzo surfaces of degree 11
Let $S$ be a Del Pezzo surface of degree $1$ defined over $\mathbb{Q}$. By [15, Theorem III.3.5] and [14, Corollary 3.54], up to a change of coordinates, we may assume that the surface $S$ has the equation
@@
$w:x:y:z]\mapsto[\lambda\cdot w+p_{3}+z\cdot p_{1}:ax+by:cx+dy:\mu\cdot z+p_{2}],$
@@
where $g_{i}=f_{i}(ax+by,cx+dy)$. This should be a multiple of the Equation 8. The degree $1$ coefficient in $w$ is $2\cdot\lambda\cdot w\cdot(p_{3}+z\cdot p_{1})$ and needs to be zero, so $p_{3}=0$ and $p_{1}=0$. Then, the degree $2$ coefficient of $z$ is $3\cdot\mu^{2}\cdot p_{2}$ so needs to be zero. This implies that every automorphism is of the form
$w:x:y:z]\mapsto[\lambda\cdot w:ax+by:cx+dy:\mu\cdot z].$
@@
where $g_{i}=f_{i}(ax+by,cx+dy)$. We find that $\lambda^{2}=\mu^{3}$, so by Lemma 6.3, there exists $u\in\mathbb{Q}^{*}$ such that $\lambda=u^{3}$ and $\mu=u^{2}$. As $w:x:y:z]=[u^{3}w:ux:uy:u^{2}z]$, we can replace $\lambda$ with $\lambda/u^{3}=1$ and $\mu$ with $\mu/u^{2}=1$. Every element is then of the form
$w:x:y:z]\mapsto[w:ax+by:cx+dy:z],$
where $\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in\operatorname{GL}_{2}(\mathbb{Q})$. Moreover the matrix is unique, as if we multiply everything with $s^{3}:s:s:s^{2}]$, then to obtain $s^{2}=s^{3}=1$ we need $s=1$. The group of automorphism is then a finite subgroup of $\operatorname{GL}_{2}(\mathbb{Q})$. In $\operatorname{GL}_{2}(\mathbb{Q})$, every finite subgroup divides $12$ (see [19, Chap. IX, section 14]), then $\mid\operatorname{Aut}(S_{\mathbb{Q}})\mid\leq 12$. Replacing again in the Equation 8, we see that $f_{i}(ax+by,cx+dy)=f_{i}$ for each $i$. The group of automorphism is then just
@@
where $\alpha^{6}=\beta^{2}=id$ and $\beta\alpha\beta=\alpha^{-1}$. Hence, the group $\operatorname{Aut}(S_{\mathbb{Q}})\cong D_{6}$ and it has $12$ elements. Moreover, If $\left\{\lambda=-{\frac{1}{5}},\mu=-{\frac{6}{5}}\right\}$, this surface $S$ contains $4$ disjoint (1)(-1)-curves, given by
@@
By contracting these four (1)(-1)-curves, there is thus a contraction morphism $\eta:S\rightarrow Y$, where $Y$ is a Del Pezzo surface of degree $5$, then $Y(\mathbb{Q})\neq\emptyset$, so the surface $S$ is rational. Indeed, this follows from [17, Theorem 3.15].